If 7.97 mol of C5H12 reacts with excess O2, how many moles of CO2 will be produced by the following combustion reaction?

C5H12 + 8O2 ⟶ 6H2O + 5CO2
1
Expert's answer
2019-10-07T06:41:55-0400

The balanced equation for considered reaction can be written as follows

"C_5H_{12} + 8 O_2 \rightarrow 6 H_2O + 5 CO_2"

"\frac{\nu(C_5H_{12})}{n(C_5H_{12})} = \frac{\nu(CO_2)}{n(CO_2)}"

where "\nu" - amount of the substance, "n" - number of the substance molecules within the balanced equation.

"\nu(CO_2)=\frac{\nu(C_5H_{12}) \cdot n(CO_2)}{n(C_5H_{12})}"

"\nu(CO_2)=\frac{7.97 \cdot 5}{1} = 39.75 \; mol"


Answer: 39.75 mol.